2. For the given equation
log base y (x-1)+ log base y ( x^2+x+1)= 1
A) solve for y.


B)Let y= f(x)and find f^-1 (x)
Thank you

Answer by stanbon(48535)

2. For the given equation
log base y (x-1)+ log base y ( x^2+x+1)= 1
A) solve for y.
= log(base y) (x-1)(x^2+x+1) = 1
y = (x-1)(x^2+x+1)
y = x^3+x^2-x^2+x-x-1 = x^3-1
----------------
B)Let y= f(x)and find f^-1 (x)
y = x^3-1
Interchange x and y to get x = y^3-1
Sove for y as follows:
y^3=x+1
y = (x+1)^(1/3)
That is f^-1(x)
================
Cheers,
Stan H.